\begin{thebibliography}{10}
\def\selectlanguageifdefined#1{
\expandafter\ifx\csname date#1\endcsname\relax
\else\language\csname l@#1\endcsname\fi}
\ifx\undefined\url\def\url#1{{\small #1}}\else\fi
\ifx\undefined\BibUrl\def\BibUrl#1{\url{#1}}\else\fi
\ifx\undefined\BibAnnote\long\def\BibAnnote#1{}\else\fi
\ifx\undefined\BibEmph\def\BibEmph#1{\emph{#1}}\else\fi

\bibitem{Ando_Itoh:1987}
\selectlanguageifdefined{english}
\BibEmph{Ando~Yu.} Calculation of transmission tunneling current across
  arbitrary potential barriers~/ Yu.~Ando, T.~Itoh~/$\!$/ J. Appl. Phys. "---
\newblock 1987. "---
\newblock Vol.~61. "---
\newblock P.~1497--1502.

\bibitem{Griffiths_Steinke:2001}
\selectlanguageifdefined{english}
\BibEmph{Griffiths~David~J.} Waves in locally periodic media~/
  David~J.~Griffiths, Carl~A.~Steinke~/$\!$/ Am. J. Phys. "---
\newblock 2001. "---
\newblock Vol.~69. "---
\newblock P.~137--154.

\bibitem{Pereyra_Castillo:2002}
\selectlanguageifdefined{english}
\BibEmph{Pereyra~Pedro}. Theory of finite periodic systems: General expressions
  and various simple and illustrative examples~/ Pedro~Pereyra,
  Edith~Castillo~/$\!$/ Phys. Rev. B. "---
\newblock 2002. "---
\newblock Vol.~65. "---
\newblock P.~205120(1)--205120(26).

\bibitem{SanchezSoto_atall:2012}
\selectlanguageifdefined{english}
The transfer matrix: A geometrical perspective~/ Luis~L.~S{\'a}nchez-Soto,
  Juan~J.~Monz{\'o}n, Alberto~G.~Barriuso, Jos{\'e}~F.~Cari{\=n}ena~/$\!$/
  Phys. Rep. "---
\newblock 2012. "---
\newblock Vol.~513. "---
\newblock P.~191--227.

\bibitem{Harwit_Harris_Kapitulnik:1986}
\selectlanguageifdefined{english}
\BibEmph{Harwit~A.~Harris~J.~S.} Calculated quasieigenstates and
  quasieigenenergies of quantum well superlattices in an applied electric
  field~/ J.~S.~Harwit, A.~Harris, A.~Kapitulnik~/$\!$/ J. Appl. Phys. "---
\newblock 1986. "---
\newblock Vol.~60. "---
\newblock P.~3211--3213.

\bibitem{Capasso_Mohammed_Cho:1986}
\selectlanguageifdefined{english}
\BibEmph{Capasso~F.} Sequential resonant tunneling through a multiquantum well
  superlattice~/ F.~Capasso, K.~Mohammed, A.~Y.~Cho~/$\!$/ Appl. Phys. Lett.
  "---
\newblock 1986. "---
\newblock Vol.~48. "---
\newblock P.~478--480.

\bibitem{Miller_etall:1985}
\selectlanguageifdefined{english}
Electric field dependence of optical absorption near the band gap of
  quantum-well structures~/ D.~A.~B.~Miller, D.~S.~Chemla, T.~C.~Damen
  [et~al.]~/$\!$/ Phys. Rev. B. "---
\newblock 1985. "---
\newblock Vol.~32, No.~2. "---
\newblock P.~1043--1060.

\bibitem{Zhou:1993}
\selectlanguageifdefined{english}
\BibEmph{Zhou~P.} Finite Difference Method~/ P.~Zhou~/$\!$/ {Numerical Analysis
  of Electromagnetic Fields. Electric Energy Systems and Engineering} {S}eries.
  "---
\newblock Heidelberg : Springer, 1993. "---
\newblock P.~427.

\bibitem{Grossmann_Roos_Stynes:2007}
\selectlanguageifdefined{english}
\BibEmph{Grossmann~C.} Finite Difference Method~/ C.~Grossmann, H.-G.~Roos,
  M.~Stynes~/$\!$/ Numerical Treatment Equations. "---
\newblock Heidelberg : Springer, 2007. "---
\newblock P.~601.

\bibitem{Kabir_Khan_Alam:1991}
\selectlanguageifdefined{english}
\BibEmph{Kabir~S. M.~F.} Application of quantum mechanical wave impedance in
  the solution of {S}chrodinger's equation in quantum wells~/ S.~M.~F.~Kabir,
  M.~R.~Khan, Alam~M.~A.~/$\!$/ Appl. Phys. Lett. "---
\newblock 1991. "---
\newblock Vol.~34, No.~12. "---
\newblock P.~1466--1468.

\bibitem{Hague_Khondker:1998}
\selectlanguageifdefined{english}
\BibEmph{Haque~A.} An efficient technique to calculate the normalized wave
  functions in arbitrary one dimensional quantum well structures~/ A.~Haque,
  A.~N.~Khondker~/$\!$/ J. Appl. Phys. "---
\newblock 1998. "---
\newblock Vol.~84, No.~10. "---
\newblock P.~5802--5804.

\bibitem{Nelin_Imamov:2010}
\selectlanguageifdefined{english}
\BibEmph{Nelin~E.~A.} Numeral {H}ilbert transform for crystal-like structures~/
  E.~A.~Nelin, V.~O.~Imamov~/$\!$/ Visn. {NTUU} {KPI} {S}er. - {R}adiotekh.
  {R}adioaparatobud. "---
\newblock 2010. "---
\newblock Vol.~41. "---
\newblock P.~80--82.

\bibitem{Babushkin_Nelin:2011_1}
\selectlanguageifdefined{english}
\BibEmph{Babushkin~A.M.} Computer simulation of quantum-size structures in the
  matlab environment~/ A.M.~Babushkin, E.A.~Nelin~/$\!$/ Visn. {NTUU} {KPI}
  {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2011. "---
\newblock Vol.~46. "---
\newblock P.~159--163.

\bibitem{Ashby:2016}
\selectlanguageifdefined{english}
\BibEmph{Ashby~A.} The study of modeled atomtronic barrier potentials through
  the impedance method : Master of science~/ {F}aculty of the {G}raduate
  {S}chool of the {U}niversity of {C}olorado. "---
\newblock Colorado, 2016. "---
\newblock 190~p.

\bibitem{Calecki_Palmier_Chomette:1984}
\selectlanguageifdefined{english}
\BibEmph{Calecki~D.~Palmier~J.~F.} Hopping conduction in multiquantum well
  structures~/ J.~F.~Calecki, D.~Palmier, A.~Chomette~/$\!$/ Journ.of Phys. C:
  Sol. Stat. Phys. "---
\newblock 1984. "---
\newblock Vol.~17, No.~28. "---
\newblock P.~5017--5030.

\bibitem{Tsu_Dohler:1975}
\selectlanguageifdefined{english}
\BibEmph{Tsu~R.} Hopping conduction in a {``}superlattice{''}~/ R.~Tsu,
  G.~D{\"{o}}hler~/$\!$/ Phys. Rev. B. "---
\newblock 1975. "---
\newblock Vol.~12, No.~2. "---
\newblock P.~680--686.

\bibitem{Lui_Fukuma:1986}
\selectlanguageifdefined{english}
\BibEmph{Lui~W.~W.} Exact solution of the {S}chrodinger equation across an
  arbitrary one-dimensional piecewise-linear potential barrier~/ W.~W.~Lui,
  M.~Fukuma~/$\!$/ Journ. of Appl. Phys. "---
\newblock 1986. "---
\newblock Vol.~60. "---
\newblock P.~1555--1559.

\bibitem{Babushkin_Nelin:2011}
\selectlanguageifdefined{english}
\BibEmph{Vodolazka~M.~V.} Resonance filtration by two-phase resonators~/
  M.~V.~Vodolazka, A.~P.~Tolstenkova, Nelin~E.~A.~/$\!$/ Visn. {NTUU} {KPI}
  {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2014. "---
\newblock Vol.~57. "---
\newblock P.~113--120.

\bibitem{Arx7:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Effective technique of numerical investigation of
  systems with complicated geometry of a potential~/ O.~I.~Hryhorchak~/$\!$/
  arXiv:2010.09372. "---
\newblock 2020. "---
\newblock P.~1--11.

\bibitem{Arx1:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Reformulation of a transmission and reflection
  problems in terms of a quantum wave impedance function~/
  O.~I.~Hryhorchak~/$\!$/ arXiv:2010.04682. "---
\newblock 2020. "---
\newblock P.~1--14.

\bibitem{Arx2:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} An application of a quantum wave impedance approach
  for solving a nonsymmetric single well problem~/ O.~I.~Hryhorchak~/$\!$/
  arXiv:2010.05583. "---
\newblock 2020. "---
\newblock P.~1--10.

\bibitem{Arx3:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Quantum wave impedance calculation for an arbitrary
  piesewise constant potential~/ O.~I.~Hryhorchak~/$\!$/ arXiv:2010.06263. "---
\newblock 2020. "---
\newblock P.~1--11.

\bibitem{Arx4:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Application of a quantum wave impedance method for
  zero-range singular potentials~/ O.~I.~Hryhorchak~/$\!$/ arXiv:2010.06930.
  "---
\newblock 2020. "---
\newblock P.~1--17.

\bibitem{Arx5:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Application of a quantum wave impedance method for
  study of infinite and semi-infinite periodic media~/ O.~I.~Hryhorchak~/$\!$/
  arXiv:2010.07632. "---
\newblock 2020. "---
\newblock P.~1--18.

\bibitem{Arx6:2020}
\selectlanguageifdefined{english}
\BibEmph{Hryhorchak~O.~I.} Numerical study of quantum mechanical systems using
  a quantum wave impedance approach~/ O.~I.~Hryhorchak~/$\!$/ arXiv:2010.08247.
  "---
\newblock 2020. "---
\newblock P.~1--18.

\bibitem{Vodolazka_Nelin:2013}
\selectlanguageifdefined{english}
\BibEmph{Vodolazka~M.~V.} Quantum{-}mechanical structures with
  delta{-}potential~/ M.~V.~Vodolazka, E.~A.~Nelin~/$\!$/ {KPI} Sci. News. "---
\newblock 2013. "---
\newblock Vol.~4. "---
\newblock P.~137--144.

\bibitem{Nelin_Liashok:2015}
\selectlanguageifdefined{english}
\BibEmph{Nelin~E.~A.} Criteria of impedance inhomogeneties approaching by
  delta-inhomogeneties~/ E.~A.~Nelin, A.~V.~Liashok~/$\!$/ Visn. {NTUU} {KPI}
  {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2015. "---
\newblock Vol.~63. "---
\newblock P.~127--135.

\bibitem{Nelin_Shulha_Zinher:2017}
\selectlanguageifdefined{english}
\BibEmph{Nelin~E.~A.} The delta-models of reactive elements and low-pass
  filters~/ E.~A.~Nelin, A.~V.~Shulha, Ya.~L.~Zinher~/$\!$/ Visn. {NTUU} {KPI}
  {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2017. "---
\newblock Vol.~69. "---
\newblock P.~72--77.

\bibitem{Nelin_Liashok:2016}
\selectlanguageifdefined{english}
\BibEmph{Nelin~E.~A.} Criteria of crystal-like structures approaching by
  impedance delta-inhomogeneities lattices~/ E.~A.~Nelin, A.~V.~Liashok~/$\!$/
  Visn. {NTUU} {KPI} {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2016. "---
\newblock Vol.~67. "---
\newblock P.~58--64.

\bibitem{Nelin_Shulha_Zinher:2018}
\selectlanguageifdefined{english}
\BibEmph{Nelin~E.~A.} Delta models of oscillatory structures and passband
  filters~/ E.~A.~Nelin, A.~V.~Shulha, Ya.~L.~Zinher~/$\!$/ Visn. {NTUU} {KPI}
  {S}er. - {R}adiotekh. {R}adioaparatobud. "---
\newblock 2018. "---
\newblock Vol.~73. "---
\newblock P.~63--68.

\end{thebibliography}

\documentclass[24pt]{article}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{amssymb,amsmath}
\usepackage{mathrsfs}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage[cp1251]{inputenc}
\usepackage{graphicx}
\usepackage{color}

\title{Analytical representation of an iterative formula for a quantum wave impedance determination in a case of a piecewise constant potential.}
\author{O. I. Hryhorchak\\
{\small Department for Theoretical Physics, Ivan Franko National
University of Lviv,}\\
{\small 12, Drahomanov Str., Lviv, UA--79005,
Ukraine}\\
\small{\it{Orest.Hryhorchak@lnu.edu.ua}}}

\def\ch{\mathop{\rm ch}\nolimits}
\def\sh{\mathop{\rm sh}\nolimits}

\begin{document}
\renewcommand{\abstractname}{Abstract}
\maketitle

\begin{abstract}
An analytical solution for a quantum wave impedance in a case of piesewise constant potential was derived. It is in fact an analytical depiction of a well-known iterative method of a quantum wave impedance determination. The expression for a transmission probability as a function of a particle energy for an arbitrary cascad of constant potentials was obtained. The application of obtained results  was illustrated on a system of double-well/barrier structures. 
\end{abstract}

\section{Introduction}

When we talk about the approximate investigation of a quantum mechanical systems with a complicated geometry of a potential the classical approach is as follows. One depicts the real potential as a cascade of constant potentials and then applies one of the various methods for getting a solution, for example a transfer matrix formalizm \cite{Ando_Itoh:1987, Griffiths_Steinke:2001, Pereyra_Castillo:2002, SanchezSoto_atall:2012,Harwit_Harris_Kapitulnik:1986,Capasso_Mohammed_Cho:1986,Miller_etall:1985},  a finite difference method \cite{Zhou:1993, Grossmann_Roos_Stynes:2007}, a quantum wave impedance approach \cite{Kabir_Khan_Alam:1991, Hague_Khondker:1998, Nelin_Imamov:2010, Babushkin_Nelin:2011_1, Ashby:2016} and others \cite{Calecki_Palmier_Chomette:1984, Tsu_Dohler:1975, Lui_Fukuma:1986, Babushkin_Nelin:2011, Nelin_Imamov:2010}. 

Generally saying we can approximate the real potential not only by a piesewise constant potential but also by a linear, parabolic, exponential forms of a potential \cite{Arx7:2020}. It is worth to say that  these methods gives us the possibility to formalize the process of quantum systems studying. But each of them has drawbacks. For example, transfer matrix formalizm demands constructing transfer matrices for each individual area of a considered potential and then performing the operation of their multiplication. And although only two matrices for each individual area of potential are required (one describes the propagation of a wave function over a region of a constant potential and the other one relates wave functions at the interfaces of different regions of a potential) it is sometimes practically too much complicated task, especially in a case when a considered potential is approximated by a large number of intervals of constant potentials.

The other very effective method of a theoretical investigation of a quantum mechanical systems is a quantum wave impedance approach \cite{Arx1:2020, Arx2:2020, Arx3:2020, Arx4:2020, Arx5:2020}.  
In \cite{Arx3:2020} we otained the well-known iterative formula for a quantum wave impedance determination and in \cite{Arx6:2020} we discussed a numerical investigation of systems with complicated geometry of a potential using this iterative approach. 

But the iterative approach also has its drawbacks. One of them is that it does not allow an explisit analysing the dependence of a quantum wave impedance on the value of a potential energy of each interval. Thus the quaestion arises, whether it is possible to depict an iterative formula for a quantum wave impedance in  an analytical form. This problem, namely obtainingan an analytical representation of  iterative formula  is the main aim of this paper. 

\section {Analytical representation of an approximate itera\-tive calculation of a quantum wave impe\-dance}

Consecutive iterative cycles of a quantum wave impedance calculation can be depicted as the multiplication of appropriate matrices. At the end we have to multiply the result matrix on the $\begin{pmatrix} 1 \\ 1 \end{pmatrix}$ vector column. The final result for a value of a quantum wave impedance we get as a fraction of a top row on a bottom one of the result vector column. We will notify such operations with the $\rightarrow$ sign. For example the following relation
\begin{eqnarray}
Z_2=z_2\frac{z_1\ch(\gamma_2l_2)-z_2\sh(\gamma_2l_2)}
{-z_1\sh(\gamma_2l_2)+z_2\ch(\gamma_2l_2)}
\end{eqnarray}
we can depict as
\begin{eqnarray}
Z_2 \rightarrow z_2
\begin{pmatrix}
z_1\ch(\gamma_2l_2) -z_2\sh(\gamma_2l_2) 
\\
-z_1\sh(\gamma_2l_2)+ z_2\ch(\gamma_2l_2) 
\end{pmatrix}
\end{eqnarray}
or
\begin{eqnarray}
Z_2 \rightarrow z_2
\begin{pmatrix}
z_1\ch(\gamma_2l_2)& -z_2\sh(\gamma_2l_2) 
\\
-z_1\sh(\gamma_2l_2)& z_2\ch(\gamma_2l_2) 
\end{pmatrix}
\begin{pmatrix}
1 
\\
1 
\end{pmatrix},
\end{eqnarray}
and vice versa
\begin{eqnarray}
z_2
\begin{pmatrix}
z_1\ch(\gamma_2l_2)& -z_2\sh(\gamma_2l_2) 
\\
-z_1\sh(\gamma_2l_2)& z_2\ch(\gamma_2l_2) 
\end{pmatrix}
\rightarrow
z_2\frac{\!z_1\ch(\gamma_2l_2)\!-\!z_2\sh(\gamma_2l_2)}
{\!-z_1\sh(\gamma_2l_2)\!+\!z_2\ch(\gamma_2l_2)}.
\end{eqnarray}
For two consecutive regions with different characteristic impedances $z_1$, $z_2$ and different widths $l_1$, $l_2$ we have:
\begin{eqnarray}
Z_2\!\!\!\!\!\! &\rightarrow& z_2
\begin{pmatrix}
z_1\ch(\gamma_2l_2)& -z_2\sh(\gamma_2l_2) 
\\
-z_1\sh(\gamma_2l_2)& z_2\ch(\gamma_2l_2) 
\end{pmatrix}
\begin{pmatrix}
z_0\ch(\gamma_1l_1)& -z_1\sh(\gamma_1l_1) 
\\
-z_0\sh(\gamma_1l_1)& z_1\ch(\gamma_1l_1) 
\end{pmatrix}
\times
\begin{pmatrix}
1 
\\
1 
\end{pmatrix}
=\nonumber\\
\!\!\!\!\!\!\!\!&=&
\begin{pmatrix}
z_0z_1\ch(\gamma_1l_1)\ch(\gamma_2l_2)-z_1^2\sh(\gamma_1l_1)\ch(\gamma_2l_2)-\\-z_1z_2\ch(\gamma_1l_1)\sh(\gamma_2l_2)+z_0z_2\sh(\gamma_1l_1)\sh(\gamma_2l_2) 
\\
\\
z_1z_2\ch(\gamma_1l_1)\ch(\gamma_2l_2)-z_0z_2\sh(\gamma_1l_1)\ch(\gamma_2l_2)-\\
-z_0z_1\ch(\gamma_1l_1)\sh(\gamma_2l_2)+z_1^2\sh(\gamma_1l_1)\sh(\gamma_2l_2) 
\end{pmatrix}\rightarrow\nonumber\\
\!\!\!\!\!\!\!\!&\rightarrow &\!\!\!\!\!\!
z_2\frac{z_0z_1\!\ch(\gamma_1l_1)\!\ch(\gamma_2l_2)\!-\!z_1^2\!\sh(\gamma_1l_1)\!\ch(\gamma_2l_2)\!-\! z_1z_2\!\ch(\gamma_1l_1)\!\sh(\gamma_2l_2)\!+\!z_0z_2\sh(\gamma_1l_1)\!\sh(\gamma_2l_2)}
{z_1z_2\!\ch(\gamma_1l_1)\!\ch(\gamma_2l_2)\!-\!z_0z_2\sh(\gamma_1l_1)\!\ch(\gamma_2l_2)\!-\! z_0z_1\!\ch(\gamma_1l_1)\!\sh(\gamma_2l_2)\!+\!z_1^2\sh(\gamma_1l_1)\!\sh(\gamma_2l_2)}.\nonumber\\
\end{eqnarray}

For $N$ consecutive regions with characteristic impedances $z_i, i=1\ldots N$ and widthes $l_i, i=1\ldots N$ we have
\begin{eqnarray}
Z_N\rightarrow z_N\prod\limits_{i=N}^1\left\{
\begin{pmatrix}
z_{i-1}\ch(\gamma_il_i)& -z_i\sh(\gamma_il_i) 
\\
-z_{i-1}\sh(\gamma_il_i)& z_i\ch(\gamma_il_i) 
\end{pmatrix}
\right\}
\begin{pmatrix}
1 
\\
1 
\end{pmatrix}.
\end{eqnarray}
Taking into account that
\begin{eqnarray}
\ch(\gamma_il_i)=\frac{e^{\gamma_il_i}+e^{-\gamma_il_i}}{2},\quad \sh(\gamma_il_i)=\frac{e^{\gamma_il_i}-e^{-\gamma_il_i}}{2}
\end{eqnarray}
we get
\begin{eqnarray}
Z_N\rightarrow\frac{z_N}{2^N}\prod\limits_{i=N}^1\left\{
e^{\gamma_il_i}
\begin{pmatrix}
z_{i-1}& -z_i 
\\
-z_{i-1}& z_i 
\end{pmatrix}
+e^{-\gamma_il_i}
\begin{pmatrix}
z_{i-1}& z_i 
\\
z_{i-1}& z_i 
\end{pmatrix}
\right\}\!\!
\begin{pmatrix}
1 
\\
1 
\end{pmatrix}\!\!.
\end{eqnarray}
After introducing following notations 
\begin{eqnarray}
Z_j^-=\begin{pmatrix}
z_{j-1}& -z_j 
\\
-z_{j-1}& z_j 
\end{pmatrix},
\qquad
Z_j^+=\begin{pmatrix}
z_{j-1}& z_j 
\\
z_{j-1}& z_j 
\end{pmatrix}
\end{eqnarray}
we finally get the  expression for the product of $N$ matrices 
\begin{eqnarray}
\prod\limits_{j=N}^1 Z_j^{i_j}=
\begin{pmatrix}
\prod\limits_{j=1}^N (z_{j-1}+i_ji_{j-1}z_j)
\\
\prod\limits_{j=1}^N{i_N}(z_{j-1}+i_ji_{j-1}z_j)
\end{pmatrix},
\end{eqnarray}
where $i_j=\pm1$.
After simple but long transformations we get the final result
\begin{eqnarray}
Z(x_N)=Z_N=z_N\frac{\sum\limits_{\{i_j\}}K(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}
{\sum\limits_{\{i_j\}}i_NK(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]},
\end{eqnarray}
where
\begin{eqnarray}
K(i_j)=\frac{1}{2^N}\prod\limits_{j=1}^N (z_{j-1}+i_ji_{j-1}z_j),
\end{eqnarray}
$z_j$ and $k_j$ are the functions of a particle energy. 
This expression can be written in the another form, namely
\begin{eqnarray}
Z(x_N)=Z_N=z_N\frac{\sum\limits_{\{i_j\}}K(i_j)\ch\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}
{\sum\limits_{\{i_j\}}i_NK(i_j)\sh\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}.
\end{eqnarray}
A condition for finding energies of bound states is as follows
\begin{eqnarray}
\frac{\sum\limits_{\{i_j\}}K(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_j\l_j\right]}
{\sum\limits_{\{i_j\}}i_NK(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}=-\frac{z_{out}}{z_N},
\end{eqnarray}
where $z_{out}$ is a characteristic impedance of the region which is on the left of a studied system.

An expression which describes the dependence of a transmission coefficient on a particle energy is as follows
\begin{eqnarray}
T(E)=1-\left|\frac{\sum\limits_{\{i_j\}}(z_N-i_Nz_0)K(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}
{\sum\limits_{\{i_j\}}(z_N+i_Nz_0)K(i_j)\exp\left[-\sum\limits_{j=1}^Ni_j\gamma_jl_j\right]}\right|^2.
\end{eqnarray}

\section{Double barrier/well systems. Analytical me\-thod}
In a paper \cite{Arx3:2020} we studied a system of two symmetric barriers/ wells using both an iterative method of a quantum wave impedance calculation and direct relations for a quantum wave impedance function. In this section we will do the same but with the help of results of the previous section.

The analytical approach which was developed in the previous section for a double barrier/well system (including nonsymmetric case) gives:
\begin{eqnarray}
Z=z_3\frac{\sum\limits_{\pm_1\pm_2\pm_3}\!\!\!\!\!
	(z_0\pm_1z_1)(z_1\pm_{1,2}z_2)(z_2\pm_{2,3}z_3)e^{\mp_1\gamma_1l_1\mp_2\gamma_2l_2\mp_3\gamma_3l_3}}
{\sum\limits_{\pm_1\pm_2\pm_3}\!\!\!\!\!\pm_3
	(z_0\pm_1z_1)(z_1\pm_{1,2}z_2)(z_2\pm_{2,3}z_3)
	e^{\mp_1\gamma_1x_1\mp_2\gamma_2x_2\mp_3\gamma_3x_3}},\nonumber\\
\end{eqnarray}
where $\pm_{1,2}=\pm_1\pm_2$, $\pm_{2,3}=\pm_2\pm_3$.
For a symmetric case we have $z_0=z_2=z$, $\gamma_0=\gamma_2=\gamma$, $z_1=z_3=\tilde{z}$, $\gamma_1=\gamma_3=\tilde{\gamma}$ and thus
\begin{eqnarray}
Z=\tilde{z}\frac{\!\!\!\!\!\sum\limits_{\pm_1\pm_2\pm_3}\!\!\!\!\!(z\pm_1\tilde{z})(\tilde{z}\pm_{1,2}z)(z\pm_{2,3}\tilde{z})e^{\mp_1\tilde{\gamma}l_1\mp_2kl_2\mp_3\tilde{\gamma}l_1}}
{\!\!\sum\limits_{\pm_1\pm_2\pm_3}\!\!\!\!\!\pm_3(z\pm_1\tilde{z})(\tilde{z}\pm_{1,2}z)(z\pm_{2,3}\tilde{z})e^{\mp_1\tilde{\gamma}l_1\mp_2kl_2\mp_3\tilde{\gamma}l_1}}\nonumber\\
\end{eqnarray}
or in the explicit form
\begin{eqnarray}
Z\!\!\!&=&\!\!\!\tilde{z}\left(\frac{}{}4\tilde{z}(\tilde{z}^2-z^2)\sh[\gamma l_2]+(\tilde{z}+z)^3e^{-\gamma l_2-2\tilde{\gamma}l_1}+(\tilde{z}-z)^3e^{-\gamma l_2+2\tilde{\gamma}l_1} -\right.\nonumber\\
&-&\left.(\tilde{z}-z)^2(\tilde{z}+z)e^{\gamma l_2-2\tilde{\gamma}l_1}-(\tilde{z}+z)^2(\tilde{z}-z)e^{\gamma l_2+2\tilde{\gamma}l_1}\right)\times\nonumber\\
&\times&\left(\frac{}{}-4z(\tilde{z}^2-z^2)\sh[\gamma l_2]+(\tilde{z}+z)^3e^{-\gamma l_2-2\tilde{\gamma}l_1}-\frac{}{}(\tilde{z}-z)^3e^{-\gamma l_2+2\tilde{\gamma}l_1}-\right.\nonumber\\
&-&(\tilde{z}-z)^2(\tilde{z}+z)e^{\gamma l_2-2\tilde{\gamma}l_1}\left.(\tilde{z}+z)^2(\tilde{z}-z)e^{\gamma l_2+2\tilde{\gamma}l_1}\right)^{-1}.
\end{eqnarray}

It is easy to transform this expression so it coincides with the expression obtained in a \cite{Arx3:2020}.

\section{Conclusions}
In this paper it was proved that the iterative process of calculation of a quantum wave impedance can be depicted analytically. So, we got the analytical solution of an equation for a quantum wave impedance with a piecewise constant potential while in  \cite{Arx3:2020} the solution was obtained in an iterative form. Analytical form of a solution is especially useful for the analysis of the relation between the geometry of a potential and an input value of a quantum wave impedance, which was demonstrated on the system of double barrier/well. The dependence of a transmission probability on a particle energy was obtained in an analytical form as well for a piesewise constant potential.

Notice, that using the analytical depiction of an iterative formula demands more computing resource for a quantum wave impedance determination that in a case of direct using iterative formula, but at the same time it gives much clear understanding of influence of each interval of a potential on the final result for a value of a quantum wave impedance. Obtained results are directly related to a design of nanodevices with the desired characteristics since they allows relating the characteristics of a device with the structure of its potential. An additional argument in favor of use a quantum wave impedance method is that it  demands ferew calculations compared to the transfer matrix approach and in many papers \cite{Kabir_Khan_Alam:1991, Hague_Khondker:1998, Vodolazka_Nelin:2013, Nelin_Liashok:2015, Nelin_Shulha_Zinher:2017, Nelin_Liashok:2016, Nelin_Shulha_Zinher:2018} it was demonstrated its efficacy for an analysis of quantum-mechanical structures with a potential which has a complicated spatial structure.

\renewcommand\baselinestretch{1.0}\selectfont


\def\name{\vspace*{-0cm}\LARGE 
	Bibliography\thispagestyle{empty}}
\addcontentsline{toc}{chapter}{Bibliography}

{\small

	\bibliographystyle{gost780u}
	\bibliography{full.bib}
	
}

\newpage

\end{document}